# DRS: Deep Question Reformulation With Structured Output

### v1

@@ 6 Experimental Results
To demonstrate the effectiveness of our proposed zero-shot DRS method, we test it on six datasets, totaling over 1,700 data points, using four different large language models. We conduct a fair and thorough comparison between our method and four baseline methods from previous research, with all accuracy scores presented in Table 3. Additionally, we adjust relevant parameters, such as temperature and the number of candidate questions, in the zero-shot DRS method and conduct multiple experiments. The results are shown in Figure 4 and Figure 3. Finally, considering GPU resources and API costs, we select a subset of test data for additional experiments with more models, as listed in Appendix B.
@@
Robust DRS. In the process of our DRS method, several steps involve using large language models to generate corresponding answers. The model’s temperature influences the output, affecting the diversity and reliability of the responses. Due to the continuity and interrelation of these steps, errors can propagate and accumulate. Therefore, we apply different temperatures to various large language models in our method, and the experimental results are presented in Figure 4.
@@
However, when the number of candidate questions is set from 2 to 5, the accuracy differences are relatively minor, typically within 3 percentage points (only Qwen2.5-7B shows a large decrease when candidate number is 4), following a general trend of initially increasing and then decreasing. Besides, the curve shows that setting the number of candidate questions to 2 or 3 yields the most ideal results, achieving high accuracy while significantly saving search time. Even more convincing is the fact that, even when the number of candidate questions is set to the least ideal value of 1, the scores obtained by our DRS method still far outperform all other baselines. This strongly demonstrates the power and remarkable robustness of our method.
@@
However, despite the slight increase in inference time, the accuracy improvements of the DRS method over the few-shot CoT method are remarkable, with increases of 111% and 160% on the two large language models, respectively. This clearly shows a trade-off between runtime and accuracy, but the accuracy gains of the DRS method far outweigh the impact of the longer runtime. Therefore, we have strong evidence to support that the DRS method is a reasonable and efficient approach.
@@ 7 Case Study
From these observations, we conclude that even advanced methods like few-shot or few-shot CoT struggle to produce satisfactory reformulations due to the lack of constraints on generated content. This lack of control often results in outputs that either overly focus on preserving the original question’s entities, leading to unanswerable questions, or prioritize answerability at the expense of entity overlap. In contrast, the DRS method effectively balances these two requirements by systematically searching for entity combinations and generating structured statements and reformulated questions containing the specified entities. This approach unlocks the full potential of large language models, enabling more effective question rephrasing for users.
@@ 8 Conclusion
In our paper, we propose a zero-shot DRS method that significantly enhances the performance of large language models on the unanswerable question reformulation task. Our method outperforms all other baselines, including the few-shot CoT approach, across six different datasets and various large language models. We also conduct extensive experiments on key parameters, such as temperature and the number of candidate questions, with results demonstrating the strong capabilities and robustness of our approach. Furthermore, our findings reveal that the previously proposed Llama2-7B evaluator lacks the capability to fairly assess the reformulated questions. We propose using GPT-4o-mini as a more reliable, accurate, and low-latency evaluator for future research. Looking ahead, we aim to develop even more effective methods to harness the full potential of large language models, helping them assist people in understanding unfamiliar documents more effectively.
@@
In this paper, our proposed zero-shot DRS method significantly enhances the ability of large language models to reformulate unanswerable questions. However, there are still the following limitations: (i) Despite the use of efficient pruning with the DFS algorithm, the process still requires multiple passes through the document, which increases computational costs. Future research could explore methods to enable large language models to complete the reformulation in a single attempt. (ii) While we conduct tests on six datasets, they do not include documents from specific disciplines, which are common in real-world applications. Future work could focus on developing datasets that cover a broader range of domains, allowing for a more comprehensive evaluation of algorithmic and large language model capabilities.
@@ Appendix A Implementation Details
Appendix B Additional Experiments
@@
Model Method    QA2 BanditQA     BBC Reddit     Yelp SQuADv2 Average GPT-4 Zero-Shot $44.0044.00$ $30.0030.00$ $28.8128.81$ $25.6625.66$ $21.5721.57$ $64.0064.00$ $35.6735.67$ Few-Shot $48.0048.00$ $29.3329.33$ $37.2937.29$ $23.0123.01$ $35.2935.29$ $60.6760.67$ $38.9338.93$ Zero-Shot DRS (ours) 78.67 72.67 71.19 66.37 62.75 70.00 70.28 Llama3.1-70B Zero-Shot $36.0036.00$ $31.3331.33$ $33.9033.90$ $19.4719.47$ $21.5821.58$ $58.0058.00$ $33.3833.38$ Few-Shot $29.3329.33$ $29.3329.33$ $27.1227.12$ $18.5818.58$ $25.4925.49$ $36.6736.67$ $27.7527.75$ Zero-Shot DRS (ours) 65.33 74.67 67.80 61.95 74.51 66.67 68.48 Gemma2-27B Zero-Shot $10.0010.00$ $12.0012.00$ $23.7323.73$ $18.5818.58$ $17.6517.65$ $42.6742.67$ $20.7720.77$ Few-Shot $28.6728.67$ $8.008.00$ $6.786.78$ $4.424.42$ $13.7313.73$ $20.6720.67$ $13.7113.71$ Zero-Shot DRS (ours) 42.00 64.00 52.54 46.02 37.25 55.33 49.52
Appendix C DRS Prompts

### v3

@@ 5 GPT Evaluator
Model Method    QA2 BanditQA     BBC    Reddit     Yelp SQuADv2 Average GPT-3.5 Zero-Shot $28.74$ $11.41$ $13.56$ $18.58$ $19.61$ $36.69$ $21.43$ Zero-Shot w/ CoT $29.15$ $15.63$ $13.56$ $20.35$ $15.69$ $43.79$ $23.03$ Few-Shot $32.79$ $22.69$ $15.25$ $23.01$ $29.41$ $38.46$ $26.94$ Few-Shot w/ CoT $44.94$ $48.78$ $16.95$ $18.58$ $25.49$ $41.22$ $32.66$ Zero-Shot DRS (ours) 81.80 73.20 62.71 75.22 66.67 62.90 70.42 GPT-4o-mini Zero-Shot $36.44$ $15.08$ $15.25$ $23.01$ $15.69$ $47.93$ $25.57$ Zero-Shot w/ CoT $49.39$ $37.50$ $42.37$ $17.70$ $35.29$ $50.10$ $38.73$ Few-Shot $36.44$ $17.80$ $16.95$ $21.24$ $21.57$ $46.55$ $26.76$ Few-Shot w/ CoT $61.13$ $52.85$ $47.46$ $43.36$ $56.86$ $59.57$ $53.54$ Zero-Shot DRS (ours) 88.26 80.16 79.66 83.19 78.43 78.30 81.33 Qwen2.5-7B Zero-Shot $46.15$ $24.05$ $25.42$ $23.89$ $25.49$ $46.70$ $31.95$ Zero-Shot w/ CoT $47.18$ $44.21$ $27.12$ $35.40$ $31.37$ $53.55$ $39.81$ Few-Shot $39.68$ $25.14$ $15.25$ $23.89$ $19.61$ $39.64$ $27.20$ Few-Shot w/ CoT $54.28$ $45.59$ $25.42$ $25.66$ $33.33$ $50.64$ $39.15$ Zero-Shot DRS (ours) 77.73 69.84 55.93 53.10 58.82 60.55 62.66 Gemma2-9B Zero-Shot $36.03$ $21.33$ $20.34$ $18.58$ $19.61$ $42.21$ $26.35$ Zero-Shot w/ CoT $29.15$ $17.12$ $22.03$ $23.01$ $19.61$ $32.94$ $23.96$ Few-Shot $26.32$ $16.17$ $25.42$ $15.04$ $17.65$ $29.39$ $21.67$ Few-Shot w/ CoT $22.67$ $18.07$ $10.17$ $24.78$ $25.49$ $27.42$ $21.43$ Zero-Shot DRS (ours) 59.92 60.73 55.93 59.29 49.02 55.62 56.75
@@ 6 Experimental Results
To demonstrate the effectiveness of our proposed zero-shot DRS method, we test it on six datasets, totaling over 1,700 data points, using four different large language models. We conduct a fair and thorough comparison between our method and four baseline methods from previous research, with all accuracy scores presented in Table 3. Additionally, we adjust relevant parameters, such as temperature and the number of candidate questions, in the zero-shot DRS method and conduct multiple experiments. The results are shown in Figure 4 and Figure 3. Finally, considering GPU resources and API costs, we select a subset of test data for additional experiments with more models, as listed in Appendix C.
@@
Robust DRS. In the process of our DRS method, several steps involve using LLMs to generate corresponding answers. The model’s temperature influences the output, affecting the diversity and reliability of the responses. Due to the continuity and interrelation of these steps, errors can propagate and accumulate. Therefore, we apply different temperatures to various LLMs in our method, and the experimental results are presented in Figure 4.
@@
However, when the number of candidate questions is set from 2 to 5, the accuracy differences are relatively minor, typically within 3 percentage points (only Qwen2.5-7B shows a larger decrease when candidate number is 4), following a general trend of initially increasing and then decreasing. Besides, the curve shows that setting the number of candidate questions to 2 or 3 yields the most ideal results, achieving high accuracy while significantly saving search time. Even more convincing is the fact that, even when the number of candidate questions is set to the least ideal value of 1, the scores obtained by our DRS method still far outperform all other baselines. This strongly demonstrates the power and remarkable robustness of our proposed method.
@@
However, despite the slight increase in inference time, the accuracy improvements of the DRS method over the few-shot CoT method are remarkable, with increases of 113% (from 32.66 to 69.55) and 160% (21.43 to 55.63) on the listed two large language models, respectively. This clearly shows a trade-off between runtime and accuracy, but the accuracy gains of the DRS method far outweigh the impact of the longer runtime. Therefore, we have strong evidence to support that the DRS method is a reasonable and efficient approach.
@@ 7 Case Study
From these observations, we conclude that even advanced methods like few-shot or few-shot CoT struggle to produce satisfactory reformulations due to the lack of constraints on generated content. This lack of control often results in outputs that either overly focus on preserving the original question’s entities, leading to unanswerable questions, or prioritize answerability at the expense of entity overlap. In contrast, the DRS method effectively balances these two requirements by systematically searching for entity combinations and generating structured statements and reformulated questions containing the specified entities. This approach unlocks the full potential of LLMs, enabling more effective question rephrasing for users.
@@ 8 Conclusion
In our paper, we propose a zero-shot DRS method that significantly enhances the performance of large language models on the unanswerable question reformulation task. Our method outperforms all other baselines, including the few-shot CoT approach, across six different datasets and various LLMs. We also conduct extensive experiments on key parameters, such as temperature and the number of candidate questions, with results demonstrating the strong capabilities and robustness of our approach. Furthermore, our findings reveal that the previously proposed Llama2-7B evaluator lacks the capability to fairly assess the reformulated questions. We propose using GPT-4o-mini as a more reliable, accurate, and low-latency evaluator for future research. Looking ahead, we aim to develop even more effective methods to harness the full potential of LLMs, helping them assist people in understanding unfamiliar documents more effectively.
@@
In this paper, our proposed zero-shot DRS method significantly enhances the ability of LLMs to reformulate unanswerable questions. However, there are still the following limitations: (i) Despite the use of efficient pruning with the DFS algorithm, the process still requires multiple passes through the document, which increases computational costs. Future research could explore methods to enable LLMs to complete the reformulation in a single attempt. (ii) While we conduct tests on six datasets, they do not include documents from specific disciplines, which are common in real-world applications. Future work could focus on developing datasets that cover a broader range of domains, allowing for a more comprehensive evaluation of algorithmic and large language model capabilities.
@@ Appendix A Implementation Details
Appendix B Large Language Models
  • GPT-3.5: A robust large language model developed by OpenAI, capable of generating text based on instructions, and highly effective across diverse natural language processing tasks.
  • GPT-4: An advanced multi-modal language model from OpenAI that accepts both image and text inputs for text generation, achieving near-human performance on various benchmarks.
  • GPT-4o-mini: A cost-efficient multi-modal model released by OpenAI on July 18, 2024, offering low latency and cost while supporting a wide range of tasks.
  • Llama3.1-70B: The 70B-parameter flagship model from the latest Llama3.1 open-source family released by MetaAI.
  • Gemmma2: The next-generation open-source model from Google, released on June 27, 2024, as an improved version of Gemmma, available in 2B, 9B, and 27B parameter configurations.
  • Qwen2.5-7B: A 7B-parameter model from Alibaba’s latest Qwen2.5 series, delivering strong performance on various benchmarks compared to other models of similar size.
Model Method    QA2 BanditQA     BBC Reddit     Yelp SQuADv2 Average GPT-4 Zero-Shot $44.00$ $30.00$ $28.81$ $25.66$ $21.57$ $64.00$ $35.67$ Few-Shot $48.00$ $29.33$ $37.29$ $23.01$ $35.29$ $60.67$ $38.93$ Zero-Shot DRS (ours) 78.67 72.67 71.19 66.37 62.75 70.00 70.28 Llama3.1-70B Zero-Shot $36.00$ $31.33$ $33.90$ $19.47$ $21.58$ $58.00$ $33.38$ Few-Shot $29.33$ $29.33$ $27.12$ $18.58$ $25.49$ $36.67$ $27.75$ Zero-Shot DRS (ours) 65.33 74.67 67.80 61.95 74.51 66.67 68.48 Gemma2-27B Zero-Shot $10.00$ $12.00$ $23.73$ $18.58$ $17.65$ $42.67$ $20.77$ Few-Shot $28.67$ $8.00$ $6.78$ $4.42$ $13.73$ $20.67$ $13.71$ Zero-Shot DRS (ours) 42.00 64.00 52.54 46.02 37.25 55.33 49.52
Appendix C Additional Experiments
@@
Appendix D DRS Prompts